OpenStudy (anonymous):

Charmaine wants to buy a shirt for $15. If the sales tax is 4% of $15,how much will she pay in sales tax?

OpenStudy (anonymous):

The sales tax is 60 cents.

OpenStudy (anonymous):

the answer i got was 1.2

OpenStudy (anonymous):

how did u do the problem

OpenStudy (anonymous):

15 times .04 equals .60

OpenStudy (anonymous):

sales tax is the percent like 4% equals .04 you multipy 15x.04

OpenStudy (nowhereman):

On the black market she will probably get it for $10 :-D

OpenStudy (anonymous):

thanks

OpenStudy (nowhereman):

Well, she wants to pay $15 including taxes. So if x is the price without taxes \[x\cdot 1.04 = 15\] and the taxes part is \[15 - x = 15 - \frac{15}{1.04}\]

OpenStudy (anonymous):

thanks